Pranks, Inc. is a manufacturer of joke and novelty products for perpetrators of practical jokes. The corporation has paid several cash dividends throughout 20Y6, the current year. It is also declaring a stock dividend to its stockholders as the calendar year-end approaches. You’ve been brought in as a consultant to assist with this process, and also to help determine whether some missing information can be determined before the distribution of the stock dividend is made. The company has two classes of stock: common stock and cumulative preferred stock.

You’ve been able to retrieve the following information so far:

Number of common shares authorized 800,000
Number of common shares issued 650,000
Par value of common shares $20
Par value of cumulative preferred shares $30
Paid-in capital in excess of par-common stock $7,000,000
Paid-in capital in excess of par-preferred stock $0
Total retained earnings before the stock dividend is declared $33,500,000
No treasury shares have been reissued.

Total Cash

Preferred Dividends

Common Dividends

Year

Dividends

Total

Per Share

Total

Per Share

20Y1 $20,000 $20,000 $0.20 $0 $0.00
20Y2 36,000 36,000 0.36 0 0.00
20Y3 79,000 34,000 0.34 45,000 0.09
20Y4 105,000 30,000 0.30 75,000 0.15
20Y5 120,000 30,000 0.30 90,000 0.18
20Y6 180,000 30,000 0.30 150,000 0.30

Cash Dividends

The accounting manager for the company prepared the schedule of cash dividends paid from 20Y1 to 20Y6 on the Pranks, Inc. panel. However, one of the reasons for Pranks, Inc.’s missing information is that the manager is away on vacation and is unreachable by phone, because he is backpacking on a remote island that does not have cell phone reception. Management would like you to determine some information from the data you’ve collected regarding its outstanding stock.

Fill in the following answers.

How many shares of common stock are outstanding? 500,000
How many shares of preferred stock are outstanding? 100,000
What is the preferred dividend as a percent of par? 1%

Points:

3 / 3

Additional Question

Shaded cells have feedback.

In which years has Pranks, Inc. paid cumulative preferred dividends in arrears? Check all that apply.

Year 1

CorrectYear 2

CorrectYear 3

Year 4

Year 5

Year 6

Points:

1 / 1

Stock Dividend

The company declared a 4% common stock dividend on December 1, and would like you to compute the following pieces of missing information. The market value of the common shares is $25.00 on December 1, and is $30.00 on the actual distribution date of the stock, December 31.

Fill in the missing information in the following table, using the information given and your work on the other panels. All “before” items are before the stock dividend was declared. All “after” items are after the stock dividend was declared and closing entries were recorded at the end of the year.

Total paid-in capital before the stock dividend $20,000.00
Total retained earnings before the stock dividend $33,500,00
Total stockholders’ equity before the stock dividend $53,500.00
Total paid-in capital after the stock dividend $??????.??
Total retained earnings after the stock dividend $??????.??
Total stockholders’ equity after the stock dividend $??????.??
Answer #1

1)

Number of Common Stock Outstanding = Total Common Stock Dividend Paid in 20Y6/Dividend Paid Per Share = 120,000/0.30 = 400000 shares

Number of Preferred Stock Outstanding = Total Preferred Dividend Paid in 2016/Dividend Paid Per Share = 30,000/0.30 = 100000 shares

Preferred Dividend as a Percent of Par = (Preferred Dividend Per Share)/Par Value*100 = 0.30/30*100 = 1%

Part 2)

The value of paid-in capital before stock dividend, total retained earnings before stock dividend and total stockholder's equity before stock dividend is given below:

Total Paid-in Capital (Common Stock) [400,000*20 + 7,000,000] 15,000,000
Total Paid-in Capital (Preferred Stock) [100,000*30] 3,000,000
Total Paid-in Capital Before Stock Dividend 20,000,000
Retained Earnings before Stock Dividend 33,500,000
Total Stockholder's Equity before Stock Dividend [20,000,000 + 33,500,000] 53,500,000

The value of paid-in capital after stock dividend, total retained earnings after stock dividend and total stockholder's equity after stock dividend is given below:

Total Paid-in Capital (Common Stock) [400,000*20 + 7,000,000 + 400,000*4%*20] 15,320,000
Total Paid-in Capital (Preferred Stock) [100,000*30] 3,000,000
Total Paid-in Capital after Stock Dividend 18,320,000
Retained Earnings after Stock Dividend [33,500,000 - 400,000*4%*25] 33,100,000
Total Stockholder's Equity before Stock Dividend [20,000,000 + 33,500,000 + (120,000 - 100,000)] 53,520,000

3)

1)

Yes, because the number of shares issued (650,000) is greater than the number of shares outstanding (400,000).

2)

The annual preferred dividend is $30,000 (100,000*0.30). The company paid a dividend of $40,000 in Year 2011. The balance amount of $10,000 (100,000*0.30 - 20,000) in arrears for 2011 is paid as follows:

Year 20Y2 = 36,000 - 30,000 = $6,000

Year 20Y3 = 34,000 - 30,000 = $4,000